Hardware changes

From PolWiki
Jump to: navigation, search

Hardware, configuration changes during Q-Weak running

November 5, 2010: First photon detector removed: (hclog entry). The second one must have been installed the same day or shortly thereafter.

November 19, 2010: Mark added 1 mm of lead in front of the photon detector: (hclog entry).

November 29, 2010: Busted laser safety box: (hclog entry).

December ??, 2010: At some poitn this month, we stopped running the signal going into the FADC through the NIM FIFO, and ran it thorugh a passive 50:50 splitter instead. I can't remember when this happened and there was no log entry.

December 7, 2010: Replaced Arshak's paddle with Moller scint. (hclog entry). I believe a second scintillator was hooked up to channel 11 at the same time.

December 8, 2010: Locked cavity in opposite laser helicity (hclog). Can't find log entry about restoration of original helicity - I'm pretty sure it was the next day.

December 9, 2010: Removed gamma-det. HV from patch (hclog entry)

January 6, 2011: Placed 2 mm lead in front of photon detector: (hclog entry).

January 13, 2011: I checked the passive splitter Mark made (elog entry). The long delay cable was removed that night and the timing adjusted (hclog entry). NOTE: it's not clear we have any useful samples once this change was done until January 26, when Wouter and Juan Carlos looked at the timing again.

January 18, 2011: Removed 1 mm lead from photon detector, now just 1 mm: (hclog entry).

February 23, 2011: Removed passive splitter - replaced with linear FAN-IN/FAN-OUT: (hclog entry).